# Implementation Study of Cost-Effective Verification for Pietrzak’s VDF in Ethereum Smart Contracts

### v4

@@ Abstract
Verifiable Delay Function (VDF) is a cryptographic concept that ensures a minimum delay before output through sequential processing, which is resistant to parallel computing. One of the significant VDF protocols academically reviewed is the VDF protocol proposed by Pietrzak. However, for the blockchain environment, the Pietrzak VDF has drawbacks including long poof size and recursive protocol computation. In this paper, we present an implementation study of Pietrzak VDF verification on Ethereum Virtual Machine (EVM). We found that the discussion in the Pietrzak’s original paper can help a clear optimization in EVM where the costs of computation are predefined as the specific amounts of gas. In our results, the cost of VDF verification can be reduced from 4M to 2M gas, and the proof length can be generated under 8 KB with the 2048-bit RSA key length, which is much smaller than the previous expectation.
@@ 1 Introduction
Our goal is to efficiently implement VDF verification within smart contracts, providing a foundation for EVM security applications utilizing VDFs. Specifically, we focus on the Pietrzak VDF [2] for two key reasons: First, the Wesolowski VDF [1], another major VDF, requires a hash-to-prime function, which is challenging to implement directly in smart contracts. Second, we hypothesize that the performance optimization technique discussed by Pietrzak [2] could significantly reduce the overall cost within the EVM environment. As a result of this focus, our research demonstrates through theoretical analysis and experiments that Pietrzak VDF verification can be effectively utilized within EVM-compatible smart contracts.
@@
  • Through theoretical analysis, we demonstrated the existence of a unique parameter that reduces the proof length. Additionally, the actual implementation results were consistent with our theoretical expectations.
  • Contrary to previous studies’ expectations, we demonstrated that the Pietrzak VDF could be utilized in the Ethereum environment with under 8 KB calldata size with the 2048-bit RSA key length.
The rest of the paper is organized as follows: Section 2 reviews related works on VDFs. Section 3 describes the target algorithm and experiment environment. In Section 4, we measure the gas costs for VDF verification on the EVM and formulate them into a function using regression. Section 5 analyzes the optimal parameters for VDF verification costs and presents experimental results that demonstrate the theoretical analysis. Finally, Section 6 concludes the paper with a discussion of further research directions. The source code for our implementation is available in the Appendix A.
@@ 2 Related Works
The concept of recursive computation for delayed output was firstly proposed as the "Time-Lock Puzzle" by Rivest et al. [8]. The Time-Lock Puzzle does not include efficient public verifiability of the solution. This was later addressed by Mahmoody et al. in their Proof of Sequential Work (PoSW) [9]. Lenstra and Wesolowski proposed Sloth, which is considered the first Verifiable Delay Function (VDF) style mechanism [3]. Boneh et al. [10] formalized the concept of VDF. Following the introduction of VDFs, various mechanisms were proposed, among which the VDFs by Wesolowski [1] and Pietrzak [2] had unique advantages in terms of efficiency and security. For the use in blockchain, Veedo and Minroot [4] were proposed for Starknet and RANDAO respectively. However, recent findings by Biryukov et al. [5] and Leurent [6] demonstrate that distributed algorithms for modular squaring exist. This implies that the fundamental assumption of using sequential squaring puzzles is not valid, highlighting the need for further research on this issue.
Not only theoretical research, but also implementation studies are crucial for approaching practical applications. However, such studies appear to have been conducted in a rather limited scope. A significant contribution is the study by Attias et al. [7], which provides a comprehensive implementation analysis of two leading VDFs, Pietrzak and Wesolowski. Their work meticulously evaluates the performance of VDFs’ evaluation, proof generation, and verification. Their work provided us with insight into the research and optimization techniques for VDFs. However, they expected the Pietrzak VDF would not be viable for blockchains. This is contrary to our results that presented the feasibility of the Pietrzak VDF on Ethereum smart contracts. Choi et al. [11] proposed three practical commit-reveal-recover protocols. They incorporated the Wesolowski VDF into Ethereum smart contracts for the purpose of demonstration. Our research covers VDF verification, which corresponds to the Proof-of-Exponentiation (PoE) stated in their work.
@@ 3 Implementation Target and Environment
In this section, we provide the target algorithm and environment. Firstly, we define the Pietrzak VDF verification algorithm we aim to implement. Secondly, we describe the experiment environment for the cost measurement.
3.1 Pietrzak VDF verification
@@
One drawback of the Pietrzak VDF is that as the time delay parameter increases, the proof size and the number of halving protocol repetitions for the verifier also increase. In the original paper proposing the Pietrzak VDF, a method to enhance efficiency was suggested. Since the halving protocol also consumes resources, the verifier can repeat the halving protocol $\tau-\delta$ times instead of $\tau$ times. Then, the verifier needs to check $x^{2^{2^{\delta}}}\stackrel{{\scriptstyle?}}{{=}}y$. Furthermore, this approach helps to reduce the size of the proof. Algorithm 2 describes this refined verification procedure.
@@ 3.2 Target Bit Length and Target Delay
In a recent study on VDF implementation using C++ [7], VDF evaluation times with the time delay parameter $T$ ranging from $2^{20}$ to $2^{25}$ can be completed within one minute with a personal CPU. The evaluation involves performing $y=x^{2^{T}}$. In other words, the target range for $\tau$ is 20 – 25 where $T=2^{\tau}$. We chose to conduct our VDF experiments within the same parameter range, as it strikes a balance between computational feasibility and practical utility while avoiding excessive resource consumption.
@@ 3.3.1 Development Environment and Tools
The development of the contract was conducted using Solidity version 0.8.26 and EVM version Cancun. At the start of the development process, it was the latest compiler version fully supported by Foundry, which is a comprehensive Ethereum development toolchain that provides a robust environment for building, testing, and debugging smart contracts. Furthermore, gas optimization is best achieved with the latest version of the compiler.
@@ 3.3.2 Optimizer Configuration
We set up Via-IR (Intermediate Representation) and the Yul optimizer in the Solidity compiler to generate efficient bytecode. Via-IR helps efficiency, security, and simplicity in code generation by deviating from the direct compilation of Solidity code to EVM bytecode. Instead, via-IR processes an intermediate step where Yul code, an intermediate representation, is generated.
In this step, we could enhance the Yul optimizer’s efficiency by replacing the msize() operation with the free memory pointer. This adjustment necessitated manual updates to the memory indexes, a process rigorously tested to ensure the integrity of our implementation. We also applied memory-safe assembly blocks in our code. This change not only aligns with Solidity’s memory model but also activates Via-IR in the compiler settings, resulting in significant gas savings.
3.3.3 Gas Measurement in Local Testing Enviornment
@@ 4.1 Data Structure
In this study, the VDFs are implemented with key sizes of 2048 and 3072 bits. However, the EVM natively supports arithmetic operators only up to 256-bit numbers. To facilitate efficient computation of numbers exceeding 256 bits, it is necessary to employ customized data formats. We revised the big number library implementation for Solidity developed by Firo [12]. The original data structure in this library includes bytes data (bytes val), bit length (uint bitlen), and a sign indicator (bool neg). Among the three elements, we deleted the sign indicator and its related low-level logic since our implementation does not require handling negative values. This library optimizes on-chain operations such as addition and subtraction by adding leading zeros for 32-byte alignment with the EVM’s memory word size, eliminating offset management. Also, it utilizes the bit length data field to optimize multiplication and comparison operations. Therefore, we format the bytes data and pre-calculate bit length off-chain before submission, significantly reducing unnecessary on-chain computations. For example, the data format for a 2048-bit representation in JSON is as shown in Listing 1. In this way, we can reduce the gas cost even though the calldata size increases. This decision is reasonable because the primary goal is to reduce the gas cost, not the transaction calldata size.
@@ 4.2 Gas Cost for Calldata and Dispatch
Considering the prediction of data and its cost, we performed experiments to test the size of calldata and the related gas usage. The related gas usage comprises of the intrinsic gas and data dispatching gas. The intrinsic gas in EVM refers the minimum cost for a transaction. It contains the constant transaction cost (21000 gas) and the cost for data supply. For verification, we need not only to allocate data in a transaction, but also to dispatch data into the verification function. The dispatching cost is proportional to the data size. Therefore, we decided to include the dispatching cost into the data cost in this study. Figure 1(a) for $\lambda=2048$ and 1(b) for $\lambda=3072$ show the calldata size with the number of proofs and the related gas usage. We can see the gas usage and the calldata size simply and linearly increase with the number of proofs.
Figure 1: Experiment results of the calldata and its dispatch cost
(a) $\lambda=2048$
(b) $\lambda=3072$
Figure 2: Experiment results of the halving algorithm cost
(a) $\lambda=2048$
(b) $\lambda=3072$
Figure 3: Experiment results of the ModExp cost
(a) $\lambda=2048$
(b) $\lambda=3072$
Finally, to describe the relationship between transaction size and gas costs, we performed a basic regression analysis. By applying the least squares method to approximate the best fit, we derive the following linear regression equations 1 and 2 for data sizes of 2048 and 3072 bytes, respectively:
@@ 4.3 Modular Exponentiation Cost
EIP-198 [14], known for "Big Integer Modular Exponentiation", introduced the ModExp (Modular Exponential) precompiled contract to the EVM, offering a streamlined and cost-efficient approach for large integer modular exponentiation.  [^2] EIP-2565 [15] contributed to further refining the gas cost model for these modular exponentiation operations by proposing adjustments to the pricing formula. As its obvious that a precompile provides the best efficinecy, we applied ModExp for both the modular exponentiation and multi-exponentiation operations in the implementation in Algorithm 2. In Figure 3(a) and Figure 3(b) illustrating the experiment results, the pricing of ModExp is linearly proportional to the number of exponentiation as $\mathcal{G}(T)=aT+b$, where we denote $\tau$ as the logarithmic value of $T$. We could get the following results from a simple regression process.

### v5

@@ Abstract
Verifiable Delay Function (VDF) is a cryptographic concept that ensures a minimum delay before output through sequential processing, which is resistant to parallel computing. One of the significant VDF protocols academically reviewed is the VDF protocol proposed by Pietrzak. However, for the blockchain environment, the Pietrzak VDF has drawbacks including long proof size and recursive protocol computation. In this paper, we present an implementation study of Pietrzak VDF verification on Ethereum Virtual Machine (EVM). We found that the discussion in the Pietrzak’s original paper can help a clear optimization in EVM where the costs of computation are predefined as the specific amounts of gas. In our results, the cost of VDF verification can be reduced from 4M to 2M gas, and the proof length can be generated under 8 KB with the 2048-bit RSA key length, which is much smaller than the previous expectation.
@@ 1 Introduction
Our goal is to efficiently implement VDF verification within smart contracts, providing a foundation for EVM security applications utilizing VDFs. Specifically, we focus on the Pietrzak VDF [2] for two key reasons: First, the Wesolowski VDF [1], another major VDF, requires a hash-to-prime function, which is challenging to implement directly in smart contracts (see B). Second, we hypothesize that the performance optimization technique discussed by Pietrzak [2] could significantly reduce the overall cost within the EVM environment. As a result of this focus, our research demonstrates through theoretical analysis and experiments that Pietrzak VDF verification can be effectively utilized within EVM-compatible smart contracts.
@@
  • Through theoretical analysis, we demonstrated the existence of a unique parameter that reduces the proof length and the execution gas fee, making the implementation more cost-efficient.
  • We demonstrated that the Pietrzak VDF could be utilized in the Ethereum environment with under 8 KB calldata size for a 2048-bit RSA key length. This calldata size represents the serialized data required for VDF verification and highlights the feasibility of implementing the Pietrzak VDF on Ethereum smart contracts, contrary to previous studies’ expectations.
The rest of the paper is organized as follows: Section 2 reviews related works on VDFs. Section 3 describes the target algorithm and experiment environment. In Section 4, we measure the gas costs for VDF verification on the EVM and formulate them into a function using regression. Section 5 analyzes the optimal parameters for VDF verification costs and presents experimental results that demonstrate the theoretical analysis. Section 6 addresses challenges in implementing VDFs on Ethereum and explores potential directions for research and optimization. Finally, Section 7 concludes the paper by summarizing our findings and outlining future research plans. The source code for our implementation is available in A.
@@ 2 Related Works
Decentralized, bias-resistant randomness is a cornerstone of blockchain security, driving extensive research into Distributed Randomness Beacon (DRB) protocols. Recent Systematization of Knowledge (SoK) studies by Raikwar and Gligoroski [8], Choi et al. [9], and Kavousi et al. [10] provide comprehensive taxonomies of DRBs, analyzing them in terms of cryptographic primitives, interactivity, unbiasability, unpredictability, and scalability. These surveys underscore the critical role of publicly verifiable randomness in preventing adversarial bias or manipulation in decentralized applications.
One cryptographic mechanism that has gained significant traction in DRB protocols is the Verifiable Delay Function (VDF). The concept of VDFs builds on earlier advances in time-delayed cryptography. Rivest et al. introduced the “Time-Lock Puzzle” [11], which envisioned recursive computation for delayed output but lacked efficient public verifiability. Mahmoody et al. later proposed Proof of Sequential Work (PoSW) [12], thereby formalizing sequential tasks in cryptographic protocols. Lenstra and Wesolowski’s Sloth [3] is often regarded as a precursor to modern VDFs, which were later formalized by Boneh et al. [13] into efficient and publicly verifiable delay primitives.
Following the initial formalization of VDFs, various constructions have emerged. Wesolowski’s [1] and Pietrzak’s [2] designs are particularly notable for their strong security proofs and computational efficiency. Beyond DRBs, VDFs have also found applications in non-interactive cryptographic timestamping [14] and DoS mitigation [15], underscoring their versatility. Pietrzak’s VDF, in particular, has been adapted in several ways: Hoffmann and Pietrzak [16] extend it to watermarkable and zero-knowledge settings, Motepalli and Jacobsen [17] apply it to blockchain bootstrapping, and it used the Pietrzak implementation by Chia Network [18]. However, these works focus primarily on conceptual or protocol-level enhancements, rather than on optimizing on-chain performance in environments like the Ethereum Virtual Machine (EVM).
In addition to these foundational VDF-based designs, other DRB solutions have been proposed that rely on various cryptographic primitives to achieve robust randomness on-chain. For example, RandRunner [19] employs trapdoor VDFs to ensure strong uniqueness, Bicorn [20] incorporates VDFs in an optimistically efficient DRB, and RandChain [21] leverages sequential Proof-of-Work to decentralize randomness generation. Raikwar [22] further examines competitive DRB protocols, highlighting different strategies for achieving resilience against adversarial manipulation. Collectively, these systems reinforce the importance of secure, verifiable primitives that balance theoretical soundness with practical feasibility in blockchain contexts.
Later proposals, such as Veedo and Minroot [4], target specific blockchain ecosystems (e.g., Starknet or RANDAO). At the same time, recent research by Biryukov et al. [5] and Leurent [6] indicates that distributed modular squaring might be possible, challenging the assumption of a strictly sequential computation. This highlights the need to reconcile theoretical security assumptions with real-world performance constraints in VDF-based DRBs.
Despite increasing interest in VDFs, implementation-focused investigations remain sparse. One of the few in-depth evaluations is by Attias et al. [7], who examined both Pietrzak’s and Wesolowski’s VDFs—covering evaluation, proof generation, and verification—and noted that large proof sizes could be problematic for blockchain environments. Choi et al. [20] later integrated Wesolowski’s VDF into Ethereum commit-reveal schemes but did not provide a cost-optimization analysis or a large-scale feasibility study.
To address this gap, our work examines the on-chain performance of Pietrzak’s VDF on Ethereum smart contracts. We concentrate on gas costs, proof sizes, and optimization strategies that render VDF-based solutions viable. By offering a comprehensive feasibility analysis for Pietrzak’s VDF in real-world blockchain environments, we extend existing DRB research and demonstrate that VDFs can be efficiently deployed within the EVM.
@@ 3 Implementation Target and Environment
In this section, we provide the target algorithm and environment. Firstly, we define the Pietrzak VDF algorithm we aim to implement. Secondly, we describe the experiment environment for the cost measurement.
3.1 Pietrzak VDF algorithms
Our research has a model with a smart contract that supports a function that performs Pietrzak VDF verification. Referring to the Pietrzak’s original paper [2], we present the evaluation procedure and the halving protocol. In the following subsections, we show how these procedures lead to proof generation and verification in a blockchain environment, culminating in a refined approach that reduces on-chain overhead in the Ethereum Virtual Machine (EVM).
3.1.1 VDF Evaluation
The Pietrzak VDF algorithm operates with the following parameters:
  • $x$: The input value for the VDF evaluation.
  • $N$: The modulus (an RSA modulus $N=p\cdot q$), where $p$ and $q$ are large primes.
  • $T$: The time delay parameter, where $T=2^{\tau}$ for some $\tau$.
  • $y$: The output value, computed as $y=x^{2^{T}}\bmod N$.
To evaluate the VDF, one exponentiates $x$ by $2^{T}\bmod N$, as shown in Equation 1:
$y\equiv x^{2^{T}}\equiv x^{2^{2^{\tau}}}\mod{N}.$ (1)
The sequential nature of repeated squaring ensures that this operation cannot be significantly parallelized, making it an effective time-delay mechanism. Algorithm 1 details this process, serving as the foundation for proving that the output $y$ can only be computed by incurring a predefined amount of real-time work. This property is crucial for applications requiring verifiable latency and resistance to parallel computation.
Algorithm 1 Evaluation of the Pietrzak VDF
1:input: $x,T,N$
2:output: $y$
3:$y\leftarrow x$
4:for $k\leftarrow 1$ to $T$ do
5:     $y\leftarrow y^{2}\mod N$
6:end for
7:return $y$
3.1.2 Pietrzak VDF proof generation
The proof generation defined in Algorithm 2 starts with the evaluation result and is used to construct the cryptographic proofs required for VDF verification. The halving protocol detailed in lines 5 – 13 optimizes the proof generation process for VDFs by reducing the verification time incrementally. At each step, the protocol halves the time parameter $T$, recalculates intermediate values $x_{i}$ and $y_{i}$, and updates these values based on a cryptographic hash function to ensure integrity. As we implement it on Ethereum, we use Keccak256 as the cryptographic hash function. This iterative reduction continues until the desired granularity is reached, ensuring that the verifier’s workload is minimized while maintaining cryptographic security. Each halving can be tracked as intermediate values $v_{i}$ are stored as part of the proof.
Algorithm 2 Pietrzak VDF proof generation
1:input: $x,T,N$
2:output: $\{\pi_{i}\}_{i=1}^{\tau}$
3:$\tau\leftarrow\lfloor\log_{2}(T)\rfloor$
4:$(x_{1},y_{1})\leftarrow(x,y)$
5:for $i\leftarrow 1$ to $\tau$ do
6:     $v_{i}\leftarrow x^{2^{T/2^{i-1}}}$
7:     $r_{i}\leftarrow keccak256(x_{i}||y_{i}||v_{i})$
8:     $x_{i+1}\leftarrow x_{i}^{r_{i}}\cdot v_{i}\mod N$
9:     if $T/2^{i-1}$ is odd then
10:         $y_{i}\leftarrow y_{i}^{2}\mod N$
11:     end if
12:     $y_{i+1}\leftarrow v_{i}^{r_{i}}\cdot y_{i}\mod N$
13:     $\pi_{i}\leftarrow v_{i}$
14:end for
15:return $\{\pi_{i}\}_{i=1}^{\tau}$
3.1.3 Pietrzak VDF verification
@@
One drawback of the Pietrzak VDF is that as the time delay parameter $T$ increases, the proof size and the number of halving protocol repetitions for the verifier also grow. To address this, Pietrzak introduced a proof-shortening parameter, $\delta$, which allows the verifier to skip $\delta$ halving rounds. Specifically, the verifier can repeat the halving protocol $\tau-\delta$ times instead of $\tau$ times, then check $x^{2^{2^{\delta}}}\stackrel{{\scriptstyle?}}{{=}}y$. The choice of $\delta$ may be determined by the system designer, the verifier, or specified by protocol parameters, depending on the implementation. This method both reduces the proof size and alleviates verification overhead. Algorithm 4 describes this refined verification procedure.
@@
Figure 1: Algorithms for Pietrzak VDF Verification and Refined Verification.
@@ 3.2 Target Bit Length and Target Delay
In a recent study on VDF implementation using C++ [7], VDF evaluation times with the time delay parameter $T=2^{\tau}$ were examined. The study found that for $\tau$ values ranging from 20 to 25, the evaluation can be completed within one minute using a personal CPU. The evaluation involves performing $y=x^{2^{T}}$. We chose to conduct our VDF experiments within the same $\tau$ range, as it strikes a balance between computational feasibility and practical utility while avoiding excessive resource consumption.
@@ 3.3.1 Development Environment and Tools
The development of the smart contract was conducted using Solidity version 0.8.26 and EVM version Cancun. At the start of the development process, it was the latest compiler version fully supported by Foundry, which is a comprehensive Ethereum development toolchain that provides a robust environment for building, testing, and debugging smart contracts. Furthermore, gas optimization is best achieved with the latest version of the compiler.
@@ 3.3.2 Optimizer Configuration
We set up viaIR (Intermediate Representation) and the Yul optimizer in the Solidity compiler to generate efficient bytecode. viaIR helps efficiency, security, and simplicity in code generation by deviating from the direct compilation of Solidity code to EVM bytecode. Instead, viaIR processes an intermediate step where Yul code, an intermediate representation, is generated. The Yul optimizer refines this intermediate representation, enhancing the IR produced by the pipeline, as well as inline assembly and utility Yul code generated by the compiler [23].
The msize() operation is incompatible with the Yul optimizer, necessitating its replacement with the free memory pointer. This adjustment required manual updates to memory indexes, which were rigorously tested to ensure the integrity of our implementation. To comply with Solidity’s memory model and enable viaIR, we incorporated memory-safe assembly blocks. These blocks ensure robust memory management and adhere to Solidity’s safety constraints, allowing the viaIR pipeline to be activated, resulting in significant gas savings [24].
3.3.3 Gas Measurement in Local Testing Environment
@@ 4.1 Data Structure
In this study, the VDFs are implemented with key sizes of 2048 and 3072 bits. However, the EVM natively supports arithmetic operators only up to 256-bit numbers. To facilitate efficient computation of numbers exceeding 256 bits, it is necessary to employ customized data formats. We revised the big number library implementation for Solidity developed by Firo [25]. The original data structure in this library includes bytes data (bytes val), bit length (uint bitlen), and a sign indicator (bool neg). Among the three elements, we deleted the sign indicator and its related low-level logic since our implementation does not require handling negative values. This library optimizes on-chain operations such as addition and subtraction by adding leading zeros for 32-byte alignment with the EVM’s memory word size, eliminating offset management. Also, it utilizes the bit length data field to optimize multiplication and comparison operations. Therefore, we format the bytes data and include the bit length alongside it, significantly reducing unnecessary on-chain computations. For example, the data format for a 2048-bit representation in JSON is as shown in Listing 1. In this way, we can reduce the gas cost even though the calldata size increases. This decision is reasonable because the primary goal is to reduce the gas cost, not the transaction calldata size.
@@ 4.2 Gas Cost for Calldata and Dispatch
To estimate the gas usage for VDF proof calldata, we predicted the data composition and its associated costs by analyzing the probability distribution of zero and non-zero bytes in the serialized data. This prediction was derived using the EVM’s gas pricing model for calldata, where zero bytes incur 4 gas units and non-zero bytes incur 16 gas units as defined in EIP-2028 [26]. Based on these predictions, we conducted experiments to measure the actual calldata size and the corresponding gas usage under various proof configurations.
The related gas usage comprises of the intrinsic gas and data dispatching gas. The intrinsic gas in EVM refers the minimum cost for a transaction. It contains the constant transaction cost (21000 gas) and the cost for data supply. For verification, we need not only to allocate data in a transaction, but also to dispatch data into the verification function. The dispatching cost is proportional to the data size. Therefore, we decided to include the dispatching cost into the data cost in this study. Figure 2(a) for $\lambda=2048$ and 2(b) for $\lambda=3072$ show the calldata size with the number of proofs and the related gas usage. We can see the gas usage and the calldata size simply and linearly increase with the number of proofs.
(a) Calldata Cost, $\lambda=2048$
(b) Calldata Cost, $\lambda=3072$
(c) Halving Cost, $\lambda=2048$
(d) Halving Cost, $\lambda=3072$
(e) ModExp Cost, $\lambda=2048$
(f) ModExp Cost, $\lambda=3072$
Figure 2: Experiment results: calldata cost, halving cost, and ModExp cost for different configurations ($\lambda=2048$ and $\lambda=3072$).
Finally, to describe the relationship between transaction size and gas costs, we performed a basic regression analysis. By applying the least squares method [27], which finds the best-fit line by minimizing the sum of the squared differences between observed and predicted values, we derive the following linear regression equations 2 and 3 for data sizes of 2048 and 3072 bytes, respectively:
@@ 4.3 Modular Exponentiation Cost
EIP-198 [28], known for "Big Integer Modular Exponentiation", introduced the ModExp (Modular Exponential) precompiled contract to the EVM, offering a streamlined and cost-efficient approach for large integer modular exponentiation.  [^2] EIP-2565 [29] contributed to further refining the gas cost model for these modular exponentiation operations by proposing adjustments to the pricing formula. As its obvious that a precompile provides the best efficinecy, we applied ModExp for both the modular exponentiation and multi-exponentiation operations in the implementation in Algorithm 4. In Figure 2(e) and Figure 2(f) illustrating the experiment results, the pricing of ModExp is linearly proportional to the number of exponentiation as $\mathcal{G}(T)=aT+b$, where we denote $\tau$ as the logarithmic value of $T$. We could get the following results from a simple regression process.